### v1

# 25mm25mm30mm30mm

Abstract.
The construction of Sweedler duality is an important tool in the theory of Hopf algebras over a field, which is a right adjoint to the dual algebra functor. In this paper, we study the Sweedler duality of Hom-algebras and their Hom-modules. We delve into the structure of Hom-coalgebras and derive the linear morphisms associated with them. Additionally, as an application, we present the (right) Hom-(co)module morphisms under the Sweedler duality.
Key words and phrases:
Sweedler duality, Hom-algebra, Hom-coalgebra, Hom-module, Hom-comodule
2020 Mathematics Subject Classification:
Primary 17A30; Secondary 17A60, 17D30
1. Introduction
Algebraic deformation theory was first studied systematically and in depth by Gerstenhaber [5] and is now one of the most important branches of algebra. In recent years, the introduction of another type of algebraic deformation algebra, namely Hom-algebra, has attracted the attention of many algebraists. In 2006, Hartwing, Larsson and Silvestrov [7] introduced Hom-Lie algebras to describe Q-transformations in Witt and Virasoro algebras. In 2008, Makhlouf and Silvestrov [10] extended the structure of Hom-Lie algebras to quasi-Hom-Lie algebras and quasi-Lie algebras where they provided a different method for constructing Hom-Lie algebras, which is based on the process of generalising the construction of Lie algebras from associative algebras by defining the Lie bracket as a commutator. Building on this foundation, they [11][15] generalized the general (co)algebraic structure to Hom-(co)algebras which essentially extends the concept of associative algebras by deforming the associativity law into an $\alpha$-associativity condition. Later, Yau [17][18][19][21] proposed the concepts of Hom-(co)modules, Hom-Hopf modules, and the duality of Hom-algebras, thus enriching the field of Hom-algebra theory. As an emerging field within algebra, (Bi)Hom-algebra is intimately connected to category theory (See,e.g.,[22][14]), combinatorics (See,e.g.,[12]), braided group representations (See,e.g.,[9]), and quantum algebras (See,e.g.,[13]), producing a wealth of research outcomes in recent years.
The exponentials of infinitesimal characters and their convolutions are known to produce characters with elegant combinatorial counterparts [2]. In a Hopf algebra, the set of characters forms a group that acquires a natural structure of an (infinite-dimensional) Lie group under specific conditions, which are commonly met in practice, such as grading. All these elements pertain to the largest dual one can consider for a Hopf algebra, known as Sweedler duality. This duality is, in a sense, connected to finite state automaton theory [3] and algebraic combinatorial theory [16].
In the finite dimensional picture, the structure of the Hom-coalgebra is obtained from the Hom-algebra by an universal dual approach, and such a conclusion is known. This naturally leads to the question that whether the above structure can also be achieved in the infinite dimensional case? In other words, is it possible to realize the induction of Hom-algebras to Hom-coalgebras (resp. Hom-modules to Hom-comodules) by constructing a duality under infinite dimensions (as we called it above the Sweedler duality)?
In this paper, we investigate the Sweedler duality of Hom-algebras and their induction of Hom-coalgebras in general. We present the new structure of the Hom-coalgebra that is induced by the Hom-algebra within the context of Sweedler duality. Moreover, we extend our findings to Hom-modules and establish the connection between Hom-modules and Hom-comodules.
The layout of this paper is as follows. In Section 2, we review some basic concepts and lemmas that discuss the relationship between Hom-algebra and Hom-coalgebra. The structure of the Hom-coalgebra under Sweedler duality is obtained in Theorem 2. A morphism of Hom-coalgebra is derived using the condition above in Corollary 2.10. In Section 3, we present a right Hom-comodule of Hom-coalgebra, which is induced by the right-module of Hom-algebra under Sweedler duality, as shown in Theorem 4. The (right) Hom-comodule morphism is also provided in Theorem 5.
Throughout this paper, for the mapping role of elements, we will note them like this $f(v):=\langle f,v\rangle$ for all $f\in G^{*}$, $v\in G$. More details about notations can be found in [5]. Throughout, we will assume that the characteristic of the field $\mathbb{K}$ is equal to zero.
2. The Sweedler duality for Hom-algebras
In this section, we will first provide a brief overview of the fundamental concepts related to the structures of Hom-(co)algebra and the interrelationship between Hom-(co)algebras (see Definition 2.3 in [20]). Subsequently, we introduce the concept of Sweedler duality, which differs from the conventional duality (see Chapter 6 in [5]).
Definition 2.1.
A Hom-algebra is a triple $(G,\mu,\alpha)$, where $G$ is a $\mathbb{K}$-linear space, $\mu:G\otimes G\rightarrow G$ is a bilinear multiplication with $\mu(g\otimes g^{\prime})=gg^{\prime}$. For simplicity when there is no confusion, we will omit the sign $\mu$. $\alpha:G\rightarrow G$ is a $\mathbb{K}$-linear space homomorphism. Otherwise it satisfies the following conditions
[…]
They can be represented by the following commutative diagrams:
${G\otimes G\otimes G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha\otimes\mu}$${G\otimes G\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu\otimes\alpha}$$\scriptstyle{\mu}$${G\otimes G}$${G,\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$               ${G\otimes G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${G\otimes G\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha\otimes\alpha}$$\scriptstyle{\mu}$${G}$${G.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$
In this context, we will refer to a Hom-algebra $(G,\mu,\alpha)$ as $G$. Let us consider two Hom-algebras, $(G,\mu,\alpha)$ and $(G^{\prime},\mu^{\prime},\alpha^{\prime})$, both defined over a field $\mathbb{K}$. Let $f:G\to G^{\prime}$ be a $\mathbb{K}$-linear map. The linear map $f$ is said to be a morphism of Hom-algebras if it satisfies the following conditions:
(2.3) $\mu^{\prime}\circ(f\otimes f)=f\circ\mu,\quad\quad f\circ\alpha=\alpha^{\prime}\circ f.$
In other words, $f$ is a morphism of Hom-algebras if it commutes with the multiplication operations and respects the Hom-structures of $G$ and $G^{\prime}$.
Definition 2.2.
Let $(H,\mu,\alpha)$ be a subspace of $G$. A Hom-subalgebra of G is defined as a subspace of G that satisfies the following conditions:
$bc\in H,\quad\alpha(b)\in H,$
for any $b,c\in H$. In particular, if $ab\in H$ and $ba\in H$, for any $a\in G$ and $b\in H$. The Hom-algebra $G$ is equipped with an ideal, which we denote by $H$.
Definition 2.3.
A Hom-coalgebra is a triple $(C,\Delta,\beta)$, where $C$ is a $\mathbb{K}$-vector space, $\Delta:C\rightarrow C\otimes C$ is a linear map with notation $\Delta(c):=\sum_{(c)}c_{(1)}\otimes c_{(2)}$ for all $c\in C$. $\beta:C\rightarrow C$ is a $\mathbb{K}$-linear space homomorphism. And it satisfies
[…]
They can be expressed by the following commutative diagrams:
${C\otimes C\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta\otimes\beta}$${C\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta\otimes\Delta}$$\scriptstyle{\Delta}$${C\otimes C}$${C,\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta}$               ${C\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta}$${C\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta\otimes\beta}$$\scriptstyle{\Delta}$${C}$${C.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$
In a similar manner, we will refer to a Hom-coalgebra $(C,\Delta,\beta)$ as $C$. A morphism of Hom-coalgebras is a linear map of the underlying $\mathbb{K}$-linear spaces that commutes with the twisting maps and the comultiplications.
Now, we will give some lemmas and describe their functions. Lemma 2.4 (see [10]) describes the dual structure of the Hom-algebra. Lemma 2.5 (see [11]) introduces the relationship between the morphism of Hom-algebras and its duality. Lemma 2.6 (see [4]) gives the quotient space properties induced by subspaces under linear spaces. Lemma 2.7 (see [11]) explain the fundamental isomorphism theorem of the Hom-algebra.
Lemma 2.4.
Let $(G,\mu,\alpha)$ is a finite-dimensional Hom-algebra, then $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra where $\mu^{*}:G^{*}\rightarrow G^{*}\otimes G^{*}$, $\alpha^{*}(a^{*})=a^{*}\alpha$.
Lemma 2.5.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two finite-dimensional Hom-algebras, then $f:G\rightarrow G^{{}^{\prime}}$ is the morphism of Hom-algebras if and only if $f^{*}:G^{{}^{\prime}*}\rightarrow G^{*}$ is the morphism of Hom-coalgebras.
Lemma 2.6.
Let $A$ and $B$ be two $\mathbb{K}$-linear spaces, $I$ and $J$ be the subspaces of $A$ and $B$ respectively. If there exists a map $\pi:A\otimes B\rightarrow A/I\otimes B/J$, we have that $\operatorname{Ker}(\pi)=A\otimes J+B\otimes I$ and $(A\otimes B)/(A\otimes J+B\otimes I)\cong A/I\otimes B/J$.
Lemma 2.7.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras over $\mathbb{K}$, $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras, $H$ is the ideal of Hom-algebra of $G$, the map $\pi:G\rightarrow G/H$ is a regular full homomorphism. If $\operatorname{Ker}\pi\subseteq\operatorname{Ker}f$, then there exists a unique morphism of Hom-algebras $g:G/H\rightarrow G^{{}^{\prime}}$, so $g\circ\pi=f$.
In order to investigate the structure of Hom-coalgebra induced by the infinite dimensional Hom-algebra and the correspondence of the finite-codimensional ideals in the morphism of Hom-algebras, we will give the definition of the Sweedler duality in Definition 2.8 and the lemma 2.9 below.
Definition 2.8.
The Sweedler duality of $G$ is based on the universal duality $G^{*}$ as follows:
$G^{\circ}:=\{f\in G^{*}\mid\operatorname{Ker}f\supseteq J,\text{ $J$ is a finite-codimensional ideal of }G\}.$
Lemma 2.9.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras over $\mathbb{K}$, $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras. If $J$ is a finite-codimensional ideal of $G^{{}^{\prime}}$, then $f^{-1}(J)$ is a finite-codimensional ideal of $G$.
Proof.
In accordance with the equation (2.3), we obtain the following which holds for any $x\in f^{-1}(J)$ and $a\in G$:
$f(x)\in J\quad\text{and}\quad f(ax)=f(a)f(x)\in J.$
Then $ax\in f^{-1}(J)$. Similarly, we can show that $xa\in f^{-1}(J)$. For any $x\in f^{-1}(J)$, it can be shown that $f(x)\in J$ and $\alpha^{\prime}(f(x))\in J$.
Observe that $f$ is a morphism of Hom-algebras. Consequently, we have
$f\circ\alpha(x)=\alpha^{\prime}\circ f(x)\in J,$
there by establishing that $\alpha(x)\in f^{-1}(J)$. We then proceed with $\alpha(f^{-1}(J))\subseteq f^{-1}(J)$. So $f^{-1}(J)$ is the ideal of Hom-algebra $G$.
We now demonstrate that $f^{-1}(J)$ is finite-codimensional. Let $g:G\to G^{\prime}/J$ be a map defined as follows: for any $a\in G$, $a=a_{1}+a_{2}$ where $a_{1}\in f^{-1}(J)$, we have $g(a)=f(a_{2})+J$. It is evident that $\operatorname{Ker}(g)=f^{-1}(J)$. Then we can induce the injection $\widetilde{g}:G/f^{-1}(J)\rightarrow G^{{}^{\prime}}/J$ ($b+f^{-1}(J)\mapsto f(b)+J$), so we have
$dim(G/f^{-1}(J))\leq dim(G^{{}^{\prime}}/J)<+\infty$
which implies that the dimension of $G/f^{-1}(J)$ is finite. Consequently, $f^{-1}(J)$ is a finite-codimensional ideal of $G$. ∎
Inspired by Liu’ construction [8] of basic algebras, it is sufficient for us to construct the structure of Hom-coalgebras via Sweedler duality now.
Theorem 1.
$G^{\circ}$ is the subspace of $G^{*}$.
Proof.
It is straightforward to verify that $G^{\circ}$ is closed under $\mathbb{K}$-multiplication. We proceed to show that $G^{\circ}$ is also closed under addition.
For any $g,h\in G^{\circ}$, let $J$ and $H$ be two finite-codimensional ideals of $G$ such that $\operatorname{Ker}g\supseteq J$ and $\operatorname{Ker}h\supseteq H$. It is clear that $J\cap H$ is an ideal of the Hom-algebra and that $\operatorname{Ker}(g+h)\supseteq J\cap H$.
The following proof explains that $J\cap H$ is finite-codimensional. Since $H/(J\cap H)\cong(J+H)/J\subseteq G/J$, the quotient $H/(J\cap H)$ is a finite-dimensional linear space. Furthermore, since $G/H\cong(G/(J\cap H))/(H/(J\cap H))$, the finite dimension of $G/H$ implies that $dim(G/(J\cap H))=dim(G/H)+dim(H/(J\cap H))$ is also finite. Consequently, $J\cap H$ is a finite-codimensional ideal of $G$. ∎
Theorem 2.
Let $(G,\mu,\alpha)$ be a Hom-algebra. Then the Sweedler duality of $G$ is a Hom-coalgebra $(G^{\circ},\Delta,\alpha^{\circ})$, where $\Delta=\mu^{*}|_{G^{\circ}}$, $\alpha^{\circ}=\alpha^{*}|_{G^{\circ}}$ and $\mu^{*}:G^{*}\rightarrow(G\otimes G)^{*}$.
Proof.
Let $f\in G^{\circ}$. We show that $\Delta(f)\in G^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(J)=0$. Then we do the formal arithmetic
$\Delta(f)(J\otimes G+G\otimes J)=f\circ\mu(J\otimes G+G\otimes J)=0,$
since $\mu(J\otimes G)\subseteq J$ and $\mu(G\otimes J)\subseteq J$. Let a map $\pi\colon G\otimes G\to G\otimes G/(J\otimes G+G\otimes J)$, then we have $\operatorname{Ker}(\pi)=J\otimes G+G\otimes J\subseteq\operatorname{Ker}(f\mu)$. By Lemma 2.7, this induces a unique morphism $\overline{f\mu}\colon G\otimes G/(J\otimes G+G\otimes J)\to\mathbb{K}$ satisfying $\overline{f\mu}\circ\pi=f\circ\mu$. That is to say that $\overline{f\mu}\in(G\otimes G/(J\otimes G+G\otimes J))^{*}$.
According to Lemma 2.6, we have $G\otimes G/(J\otimes G+G\otimes J)\cong(G/J)\otimes(G/J)$ and $G/J$ is finite, it is easy to see that
$\overline{f\mu}\in(G/J)^{*}\otimes(G/J)^{*}.$
Let $\overline{f\mu}=\sum_{i,j=1}^{n}k_{ij}\overline{e_{i}}^{*}\otimes\overline{e_{j}}^{*}$, where $\overline{e_{i}}^{*}\in(G/J)^{*}$ is the duality base elements corresponding to a canonical map $\tilde{\pi}\colon G\to G/J$. So for any $a,b\in G$, we have
$\langle f\circ\mu,a\otimes b\rangle$ $=\langle\overline{f\mu}\circ\pi,a\otimes b\rangle$
$=\langle\overline{f\mu},\pi(a\otimes b)\rangle$
$=\langle\overline{f\mu},\tilde{\pi}(a)\otimes\tilde{\pi}(b)\rangle$
$=\sum_{i,j=1}^{n}k_{ij}\langle\overline{e}_{i}^{*},\tilde{\pi}(a)\rangle\langle\overline{e}_{j}^{*},\tilde{\pi}(b)\rangle.$
Denote $e_{i}^{*}=\overline{e}_{i}^{*}\circ\tilde{\pi}$ Since $\operatorname{Ker}(e_{i}^{*})\supseteq J$, we have $e_{i}^{*}\in G^{\circ}$. Then it is straightforward to see
$\Delta(f)=f\circ\mu=\sum\limits_{i,j=1}^{n}k_{ij}e_{i}^{*}\otimes e_{j}^{*}.$
Therefore, we conclude that $\Delta(f)\in G^{\circ}\otimes G^{\circ}$.
For any $g\in G^{\circ}$, there exists a finite-codimensional ideal $J$ of $G$ such that $\operatorname{Ker}(g)\supseteq J$. Then, since $\alpha(J)\subseteq J$ and $g(J)=0$, we have
$\langle\alpha^{\circ}(g),J\rangle=\langle\alpha^{*}(g),J\rangle=\langle g,\alpha(J)\rangle=0.$
We therefore have $\operatorname{Ker}(\alpha^{\circ}(g))\supseteq J$, which implies that $\alpha^{\circ}(g)\in G^{\circ}$. So we prove that $\alpha^{\circ}(G^{\circ})\subseteq G^{\circ}$.
Then we will prove that $\Delta\circ\alpha^{\circ}=(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta$. Using the equation (2.2), for any $c^{*}\in G^{\circ}$ and $x,y\in G$,
$\langle\Delta\circ\alpha^{\circ}(c^{*}),x\otimes y\rangle$ $=\langle c^{*},\alpha\circ\mu(x\otimes y)\rangle$
$=\langle c^{*},\mu\circ(\alpha\otimes\alpha)(x\otimes y)\rangle$
$=\langle\Delta(c^{*}),\alpha(x)\otimes\alpha(y)\rangle$
$=\sum_{(c^{*})}\langle c^{*}_{(1)}\otimes c^{*}_{(2)},\alpha(x)\otimes\alpha(y)\rangle$
$=\sum_{(c^{*})}\langle\alpha^{\circ}(c^{*}_{(1)}),x\rangle\langle\alpha^{\circ}(c^{*}_{(2)}),y\rangle$
$=\langle(\alpha^{\circ}\otimes\alpha^{\circ})(\sum_{(c^{*})}c^{*}_{(1)}\otimes c^{*}_{(2)}),x\otimes y\rangle$
$=\langle(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta(c^{*}),x\otimes y\rangle.$
Therefore we have $\Delta\circ\alpha^{\circ}=(\alpha^{\circ}\otimes\alpha^{\circ})\circ\Delta$.
It is sufficient to verify that $(\alpha^{\circ}\otimes\Delta)\circ\Delta=(\Delta\otimes\alpha^{\circ})\circ\Delta.$ For any $c^{*}\in G^{\circ}$ and $x,y,z\in G$, we have
$\langle(\alpha^{\circ}\otimes\Delta)\circ\Delta(c^{*}),x\otimes y\otimes z\rangle$ $=\langle(\alpha^{\circ}\otimes\Delta)\circ(\sum_{(c^{*})}c^{*}_{(1)}\otimes c^{*}_{(2)}),x\otimes y\otimes z\rangle$
$=\sum_{(c^{*})}\langle\alpha^{\circ}(c^{*}_{(1)})\otimes\Delta(c^{*}_{(2)}),x\otimes y\otimes z\rangle$
$=\sum_{(c^{*})}\langle\alpha^{\circ}(c^{*}_{(1)}),x\rangle\langle\Delta(c^{*}_{(2)}),y\otimes z\rangle$
$=\sum_{(c^{*})}\langle c^{*}_{(1)},\alpha(x)\rangle\langle c^{*}_{(2)},yz\rangle$
$=\sum_{(c^{*})}\langle c^{*}_{(1)}\otimes c^{*}_{(2)},\alpha(x)\otimes yz\rangle$
$=\langle\Delta(c^{*}),\alpha(x)\otimes yz\rangle$
$=\langle c^{*},\alpha(x)yz\rangle$
Similarly, we can get
$\langle(\Delta\otimes\alpha^{\circ})\circ\Delta(c^{*}),x\otimes y\otimes z\rangle$ $=\langle c^{*},xy\alpha(z)\rangle.$
Due to the Hom-associativity (2.1) of $G$, we have $\alpha(x)yz=xy\alpha(z)$. Thus $(\alpha^{\circ}\otimes\Delta)\circ\Delta=(\Delta\otimes\alpha^{\circ})\circ\Delta$. Therefore we prove this theorem. ∎
In the absence of specific instructions, we will examine the properties of Hom-coalgebra $(G^{\circ},\Delta,\alpha^{\circ})$, the Sweedler duality of $G$ as outlined in Theorem 2. In light of the aforementioned considerations, we proceed to derive the morphism of Hom-coalgebras as follows.
Corollary 2.10.
Let $(G,\mu,\alpha)$ and $(G^{{}^{\prime}},\mu^{{}^{\prime}},\alpha^{{}^{\prime}})$ be two Hom-algebras, $(G^{\circ},\Delta,\alpha^{\circ})$ and $(G^{\prime\circ},\Delta^{\prime},\alpha^{\prime\circ})$ two corresponding Sweedler dualities. The map $f:G\rightarrow G^{{}^{\prime}}$ is a morphism of Hom-algebras. Note that the map $f^{\circ}=f^{*}|_{G^{\prime\circ}}:G^{\prime\circ}\rightarrow G^{*}$. Thus, $f^{\circ}$ is a morphism of Hom-coalgebras.
Proof.
Our primary objective is to demonstrate that $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$. For any $b^{*}\in G^{\prime\circ}$, since $J$ is a finite-codimensional ideal of $G^{\prime}$ and is included in $\operatorname{Ker}(b^{*})$, we can utilize Lemma 2.9 to deduce that $f^{-1}(J)$ constitutes a finite-codimensional ideal of the Hom-algebra $G$.
Notice that
$\langle f^{\circ}(b^{*}),f^{-1}(J)\rangle$ $=\langle b^{*},f\circ f^{-1}(J)\rangle$
$=\langle b^{*},J\rangle$
$=0,$
which shows that $\operatorname{Ker}(f^{\circ}(b^{*}))\supseteq f^{-1}(J)$. In other words, $f^{\circ}(b^{*})\in G^{\circ}$. Consequently, it can be concluded that $f^{\circ}(G^{\prime\circ})\subseteq G^{\circ}$.
Using the equation (2.3), for any $b^{*}\in G^{\prime\circ}$ and $a,c\in G^{\circ}$, we have
$\langle\Delta\circ f^{\circ}(b^{*}),a\otimes c\rangle$ $=\langle b^{*},f\circ\mu(a\otimes c)\rangle$
$=\langle b^{*},\mu^{\prime}\circ(f\otimes f)(a\otimes c)\rangle$
$=\langle\Delta^{\prime}(b^{*}),f(a)\otimes f(c)\rangle$
$=\sum_{(b^{*})}\langle b^{*}_{(1)}\otimes b^{*}_{(2)},f(a)\otimes f(c)\rangle$
$=\sum_{(b^{*})}\langle b^{*}_{(1)},f(a)\rangle\langle b^{*}_{(2)},f(c)\rangle$
$=\sum_{(b^{*})}\langle f^{\circ}(b^{*}_{(1)}),a\rangle\langle f^{\circ}(b^{*}_{(2)}),c\rangle$
$=\langle(f^{\circ}\otimes f^{\circ})(\sum_{(b^{*})}b^{*}_{(1)}\otimes b^{*}_{(2)}),a\otimes c\rangle$
$=\langle(f^{\circ}\otimes f^{\circ})\circ\Delta^{\prime}(b^{*}),a\otimes c\rangle.$
Thus, we obtain $\Delta\circ f^{\circ}=(f^{\circ}\otimes f^{\circ})\circ\Delta^{\prime}$.
For any $b^{*}\in G^{\prime\circ}$ and $a\in G$, we have
$\langle\alpha^{\circ}\circ f^{\circ}(b^{*}),a\rangle$ $=\langle b^{*},f\circ\alpha(a)\rangle$
$=\langle b^{*},\alpha^{{}^{\prime}}\circ f(a)\rangle$
$=\langle f^{\circ}\circ\alpha^{\prime\circ}(b^{*}),a\rangle.$
Therefore, it follows that $\alpha^{\circ}\circ f^{\circ}=f^{\circ}\circ\alpha^{\prime\circ}$, which is the desired result. ∎
3. The Sweedler duality for Hom-module
In this section, we present the structure of Hom-(co)modules and the morphisms of Hom-(co)modules (see Definition 4.1 in [20]). Subsequently, the Sweedler duality of Hom-module is discussed (see [8]).
Definition 3.1.
Let $(G,\mu,\alpha)$ be a Hom-algebra. A right Hom-module of $G$ is a triple $(M,\psi,\gamma)$ where $M$ is a $\mathbb{K}$-linear space, a map $\psi:M\otimes G\rightarrow M$ with $\psi(m\otimes g)=m\cdot g$ and a map $\gamma:M\rightarrow M$ satisfying:
[…]
We will refer to a right Hom-module $(M,\psi,\gamma)$ as $M$. Let $(M,\psi,\gamma)$ and $(N,\tilde{\psi},\tilde{\gamma})$ be two right Hom-modules of $G$. A map $\sigma:M\rightarrow N$ is said to be a morphism of right Hom-modules when the following conditions hold:
(3.1) $\sigma\circ\psi=\tilde{\psi}\circ(\sigma\otimes\alpha),\quad\quad\tilde{\gamma}\circ\sigma=\sigma\circ\gamma.$
Definition 3.2.
Let $(C,\Delta,\beta)$ be a Hom-coalgebra. A right Hom-comodule of $C$ is a triple $(A,\phi,\epsilon)$ where $A$ is a $\mathbb{K}$-linear space, a map $\phi:A\rightarrow A\otimes C$ with $\psi(a)=\sum_{(a)}a_{(0)}\otimes a_{(1)}$ and a map $\gamma:M\rightarrow M$ satisfying:
[…]
We will refer to a right Hom-comodule $(A,\phi,\epsilon)$ as $A$. Let $(A,\phi,\epsilon)$ and $(A^{\prime},\phi^{\prime},\epsilon^{\prime})$ be two right Hom-comodule of $C$. A map $f:A\rightarrow A^{\prime}$ is called a morphism of right Hom-comodules if it satisfies:
$\phi^{\prime}\circ f=(f\otimes\beta)\circ\phi,\quad\quad\epsilon^{\prime}\circ f=f\circ\epsilon.$
Definition 3.3.
The Sweedler duality of the right Hom-module $M$ based on the universal duality $M^{*}$ is defined by
[…]
Then we will give the isomorphism in dual spaces (see [8]) and the structure of Hom-comodule induced by the universal duality as follows.
Lemma 3.4.
Let $U$ and $V$ be two $\mathbb{K}$-linear space. Denote $\rho:U^{*}\otimes V^{*}\rightarrow(U\otimes V)^{*}$ with corresponding rule $\rho(f\otimes g)(x\otimes y)=f(x)g(y)$, for any $f\in U^{*}$, $g\in V^{*}$, $x\in U$ and $y\in V$. If $\operatorname{dim}U<+\infty$ or $\operatorname{dim}V<+\infty$, $\rho$ is the isomorphism of the linear spaces.
Lemma 3.5.
Let $(G,\mu,\alpha)$ be a finite-dimensional Hom-algebra and $(M,\psi,\gamma)$ be a right Hom-module of $G$. Then, $(M^{*},\psi^{*},\gamma^{*})$ is the right Hom-comodule of the Hom-coalgebra $(G^{*},\mu^{*},\alpha^{*})$, where $\psi^{*}:M^{*}\rightarrow M^{*}\otimes G^{*}$ and $\gamma^{*}:M^{*}\rightarrow M^{*}$.
Proof.
By applying Lemma 2.4, we can get $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra. Then we want to show that $\psi^{*}\circ\gamma^{*}=(\gamma^{*}\otimes\alpha^{*})\circ\psi^{*}$. For any $m^{*}\in M$, $m\in M$ and $g\in G$, we have
$\langle\psi^{*}\circ\gamma^{*}(m^{*}),m\otimes g\rangle$ $=\langle m^{*},\gamma\circ\psi(m\otimes g)\rangle$
$=\langle m^{*},\psi\circ(\gamma\otimes\alpha)(m\otimes g)\rangle$
$=\langle\psi^{*}(m^{*}),\gamma(m)\otimes\alpha(g)\rangle$
$=\sum_{(m^{*})}\langle m^{*}_{(0)}\otimes m^{*}_{(1)},\gamma(m)\otimes\alpha(g)\rangle$
$=\sum_{(m^{*})}\langle\gamma^{*}(m^{*}_{(0)}),m\rangle\langle\alpha^{*}(m^{*}_{(1)}),g\rangle$
$=\sum_{(m^{*})}\langle(\gamma^{*}\otimes\alpha^{*})(m^{*}_{(0)}\otimes m^{*}_{(1)}),m\otimes g\rangle$
$=\langle(\gamma^{*}\otimes\alpha^{*})\circ\psi^{*}(m^{*}),m\otimes g\rangle.$
Therefore we have $\psi^{*}\circ\gamma^{*}=(\gamma^{*}\otimes\alpha^{*})\circ\psi^{*}$.
Then it is sufficient for us to verify that $(\psi^{*}\otimes\alpha^{*})\circ\psi^{*}=(\gamma^{*}\otimes\mu^{*})\circ\psi^{*}$. For any $a^{*}\in M^{*}$ and $x\in M,y,z\in G$,
$\langle(\psi^{*}\otimes\alpha^{*})\circ\psi^{*}(a^{*}),x\otimes y\otimes z\rangle$ $=\langle(\psi^{*}\otimes\alpha^{*})(\sum_{(a^{*})}a^{*}_{(0)}\otimes a^{*}_{(1)}),x\otimes y\otimes z\rangle$
$=\sum_{(a^{*})}\langle\psi^{*}(a^{*}_{(0)}),x\otimes y\rangle\langle\alpha^{*}(a^{*}_{(1)}),z\rangle$
$=\sum_{(a^{*})}\langle a^{*}_{(0)},\psi(x\otimes y)\rangle\langle a^{*}_{(1)},\alpha(z)\rangle$
$=\langle\psi^{*}(a^{*}),(\psi\otimes\alpha)(x\otimes y\otimes z)\rangle$
$=\langle a^{*},\psi\circ(\psi\otimes\alpha)(x\otimes y\otimes z)\rangle.$
Similarly, we have $\langle(\gamma^{*}\otimes\mu^{*})\circ\psi^{*}(a^{*}),x\otimes y\otimes z\rangle=\langle a^{*},\psi\circ(\gamma\otimes\mu)(x\otimes y\otimes z)\rangle$. According to the equation (3.1), It can be shown that $(\psi^{*}\otimes\alpha^{*})\circ\psi^{*}=(\gamma^{*}\otimes\mu^{*})\circ\psi^{*}$. This completes the proof of the lemma. ∎
Now, we will construct the structure of Hom-comodule via the Sweedler duality.
Theorem 3.
$M^{\circ}$ is the subspace of $M^{*}$.
Proof.
Firstly, it can be observed that $M^{\circ}$ is closed under $\mathbb{K}$-multiplication. So it is sufficient for us to show that $M^{\circ}$ is closed under addition.
Let $f,g\in M^{\circ}$, there exist finite-codimensional ideals $I$ and $J$ of $G$ such that $f(M\cdot I)=0$ and $g(M\cdot J)=0$. By Theorem 1, we can show that $I\cap J$ is an ideal of $G$ and $G/(I\cap J)$ is finite-dimensional. Then, we have $(f+g)(M\cdot(I\cap J))=0$. And we may conclude that $f+g\in M^{\circ}$, thereby completing the proof. ∎
Theorem 4.
Let $(M,\psi,\gamma)$ be a right Hom-module of a Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, then its Sweedler duality $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ is a right Hom-comodule of $G^{\circ}$ where $\psi^{\circ}=\psi^{*}|_{M^{\circ}}$, $\gamma^{\circ}=\gamma^{*}|_{M^{\circ}}$ and $\psi^{*}:M^{*}\rightarrow(M\otimes G)^{*}$.
Proof.
Noticing $f\in M^{\circ}$, we verify that $\psi^{\circ}(f)\in M^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(M\cdot J)=0$. Since $\alpha$ is the automorphism map of $G$, $\gamma(M)\subseteq M$ and $JG\subseteq J$, we have
$\psi^{\circ}(f)(M\cdot J\otimes G+M\otimes J)$ $=\psi^{\circ}(f)(M\cdot J\otimes\alpha(G)+M\otimes J)$
$=f\circ\psi(M\cdot J\otimes\alpha(G)+M\otimes J)$
$=f(\psi(M\cdot J\otimes\alpha(G))+\psi(M\otimes J))$
$=f(\psi(\gamma(M)\otimes JG)+\psi(M\otimes J))$
$=0.$
Then we denote $\pi:M\otimes G\rightarrow M\otimes G/(M\cdot J\otimes G+M\otimes J)$ and $\operatorname{Ker}(\pi)=M\cdot J\otimes G+M\otimes J\subseteq\operatorname{Ker}(f\circ\psi).$
By Lemma 2.6, we can induce a unique morphism $\overline{f\psi}:M\otimes G/(M\cdot J\otimes G+M\otimes J)\rightarrow\mathbb{K}$ satisfying $\overline{f\psi}\circ\pi=f\circ\psi$, that is to say
$\overline{f\psi}\in\left(M\otimes G/(M\cdot J\otimes G+M\otimes J)\right)^{*}.$
According to the Lemma 3.4, we can get $M\otimes G/(M\cdot J\otimes G+M\otimes J)\cong(M/M\cdot J)\otimes(G/J)$ with finite-dimensional linear space $G/J$, so we have
$\overline{f\psi}\in(M/M\cdot J)^{*}\otimes(G/J)^{*}.$
Denote two canonical maps $\pi_{1}:M\rightarrow M/M\cdot J$ and $\pi_{2}:G\rightarrow G/J$. And then we denote $\overline{f\psi}=\sum^{n}_{i,j=1}k_{ij}\overline{d_{i}}^{*}\otimes\overline{e_{j}}^{*}$ where $\overline{d_{i}}^{*}\in(M/M\cdot J)^{*}$, $\overline{e_{j}}^{*}\in(G/J)^{*}$ are the dual base elements. After a similar discussion as in Theorem 2, for any $x\in M$ and $y\in G$, we can get that
$\langle f\circ\psi,x\otimes y\rangle=\sum_{i,j=1}^{n}k_{ij}\langle\overline{d_{i}}^{*},\pi_{1}(x)\rangle\langle\overline{e_{j}}^{*},\pi_{2}(y)\rangle.$
Denote $d_{i}^{*}=\overline{d_{i}}^{*}\circ\pi_{1}$ and $e_{j}^{*}=\overline{e_{j}}^{*}\circ\pi_{2}$. Since $d_{i}^{*}(M\cdot J)=0$ and $\operatorname{Ker}(e_{j}^{*})\supseteq J$, we have $d_{i}^{*}\in M^{\circ}$ and $e_{j}^{*}\in G^{\circ}$. Thus we can show that
$\psi^{\circ}(f)=f\circ\psi=\sum_{i,j=1}^{n}k_{ij}d_{i}^{*}\otimes e_{j}^{*}.$
Consequently, $\psi^{\circ}(f)\in M^{\circ}\otimes G^{\circ}$.
For any $m\in M^{\circ}$, there exists a finite-codimensional ideal $J$ of $G$ such that $m(M\cdot J)=0$. Then we have
$\langle\gamma^{\circ}(m),M\cdot J\rangle=\langle m,\gamma(M\cdot J)\rangle=0,$
since $\gamma(M\cdot J)=\gamma(M)\cdot\alpha(J)\subseteq M\cdot J$. Then we have $\gamma^{\circ}(m)(M\cdot J)=0$, which implies $\gamma^{\circ}(m)\in M^{\circ}$. Thus, $\gamma^{\circ}(M^{\circ})\subseteq M^{\circ}$. Then by Lemma 3.5 and the relationship between the Sweedler duality and the universal duality we can show that $\psi^{\circ}\circ\gamma^{\circ}=(\gamma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}$ and $(\psi^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}=(\gamma^{\circ}\otimes\Delta)\circ\psi^{\circ}$. Hence, $M^{\circ}$ is the right Hom-comodule of $G^{\circ}$. ∎
In the subsequent proof, we will study the properties of Hom-comodule $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ proposed in Theorem 4. Based on the above considerations, we continue to derive the morphism of Hom-comodules as follows.
Theorem 5.
Let $(M,\psi,\gamma)$ and $(N,\psi^{\prime},\gamma^{\prime})$ be two right Hom-modules over Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ and $(N^{\circ},\psi^{\prime\circ},\gamma^{\prime\circ})$ be two corresponding Sweedler dualities over $G^{\circ}$. Let $\sigma:M\rightarrow N$ be a morphism of right Hom-modules. Note that $\sigma^{\circ}:=\sigma^{*}|_{N^{\circ}}:N^{\circ}\rightarrow M^{*}$. Then $\sigma^{\circ}$ is a morphism of Hom-comodules over $G^{\circ}$
Proof.
For all $f\in N^{\circ}$, let $J$ be a finite-codimensional ideal of $G$ such that $f(N\cdot J)=0$. By the properties of the morphism of the right Hom-modules, we have
$\sigma^{\circ}(f)(M\cdot J)$ $=\sigma^{\circ}(f)\circ\psi(M\otimes J)$
$=f(\sigma\circ\psi(M\otimes J))$
$=f(\psi^{\prime}\circ(\sigma\otimes\alpha)(M\otimes J))$
$=f(\sigma(M)\cdot\alpha(J))$
$=0,$
since $\sigma(M)\subseteq N$ and $\alpha(J)\subseteq J$. Hence $\sigma^{\circ}(N^{\circ})\subseteq M^{\circ}$. Then we will prove that $(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}=\psi^{\circ}\circ\sigma^{\circ}$. For any $b^{*}\in N^{\circ}$ and $x\in M,y\in G$, we have
$\langle(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}(b^{*}),x\otimes y\rangle$ $=\sum_{(b^{*})}\langle(\sigma^{\circ}\otimes\alpha^{\circ})(b^{*}_{(0)}\otimes b^{*}_{(1)}),x\otimes y\rangle$
$=\sum_{(b^{*})}\langle\sigma^{\circ}(b^{*}_{(0)}),x\rangle\langle\alpha^{\circ}(b^{*}_{(1)}),y\rangle$
$=\langle\psi^{\prime\circ}(b^{*}),(\sigma\otimes\alpha)(x\otimes y)\rangle$
$=\langle b^{*},\psi^{\prime}\circ(\sigma\otimes\alpha)(x\otimes y)\rangle$
$=\langle b^{*},\sigma\circ\psi(x\otimes y)\rangle$
$=\langle\psi^{\circ}\circ\sigma^{\circ}(b^{*}),x\otimes y\rangle.$
Thus we obtain $(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}=\psi^{\circ}\circ\sigma^{\circ}$.
Then we vertify $\gamma^{\circ}\circ\sigma^{\circ}=\sigma^{\circ}\circ\gamma^{\prime\circ}$. For all $c^{*}\in N^{\circ}$ and $x\in M$,
$\langle\gamma^{\circ}\circ\sigma^{\circ}(c^{*}),x\rangle$ $=\langle c^{*},\sigma\circ\gamma(x)\rangle$
$=\langle c^{*},\gamma^{\prime}\circ\sigma(x)\rangle$
$=\langle\sigma^{\circ}\circ\gamma^{\prime\circ}(c^{*}),x\rangle.$
Thus it follows that $\gamma^{\circ}\circ\sigma^{\circ}=\sigma^{\circ}\circ\gamma^{\prime\circ}$. Therefore this theorem is proved. ∎
In the special case when $M=G$, we have $M\cdot J=GJ\subseteq J$. Then the Sweedler duality of the Hom-module $M$ reduces to the general Sweedler duality of the Hom-algebra.
Acknowledgements S. Wang was partially supported by NSFC grant No. 12271089, and C. Zhang was partially supported by NSFC grant No. 12101111.